OpenStudy (anonymous):

Find the vertex of f(X)= 3x^-gx+8

OpenStudy (anonymous):

i guess u mean x^2-gx+8?

OpenStudy (dumbcow):

x= -b/2a = g/6 y= 3(g/6)^2 -g(g/6) +8 y = g^2/12 -g^2/6 +8 y = -g^2/12 +8 vertex: (g/6, -g^2/12 +8)